The age of two persons differ by 20 years. If 5 years ago, the elder one be 5 times as old as the younger one.what is their present age?
Answer by dfrazzetto(283)

e = elder
y = younger
e - y = 20
e = 20 + y
e - 5 = 5(y-5)
20 + y - 5 = 5y - 25
y + 15 = 5y - 25
4y = 40
y = 10
e = 30
CHECK:
30 - 5 = 5(10-5)
25 = 5*5= 25